You won the lottery and plan to buy your parents the coastal home they've always wanted. You pick out a $500,000 villa on the water, and make a $60,000 down-payment in cash. You take out a 30 year mortgage of $440,000 at 7.7% per year, and begin making the monthly mortgage payments.

After two years, your parents decide the house is too big and want to move.  

a) What is the unpaid balance on the mortgage after making two years of payments? (In other words, how much do you still owe on the house?)

b) How much total interest did you pay over two years?
Answer #1

a)

monthly payment

=440000/((1-(1+(7.7%/12))^(-30*12))/(7.7%/12))

=3137.025

What is the unpaid balance on the mortgage after making two years of payments

=3137.025*((1-(1+(7.7%/12))^(-28*12))/(7.7%/12))

=431888.85

b)

How much total interest did you pay over two years

=3137.025*24-(440000-431888.85)

=67177.45

the above is answer..